# Quasiprobability Thermodynamic Uncertainty Relation

### v1

@@ Abstract
We derive a quantum extension of the thermodynamic uncertainty relation where dynamical fluctuations are quantified by the Terletsky–Margenau–Hill quasiprobability, a quantum generalization of the classical joint probability. The obtained inequality plays a complementary role to existing quantum thermodynamic uncertainty relations, focusing on observables’ change rather than exchange of charges through jumps and respecting initial coherence. Quasiprobabilities show anomalous behaviors that are forbidden in classical systems, such as negativity; we reveal that such behaviors are necessary to reduce dissipation beyond classical limitations and show that they are stronger requirements than that the state has quantum coherence. To illustrate these statements, we employ a model that can exhibit a dissipationless heat current, which would be prohibited in classical systems; we construct a state that has much coherence but does not lead to a dissipationless current due to the absence of anomalous behaviors in quasiprobabilities.
@@
While extending the TUR to the quantum regime is a crucial problem gathering much attention recently [4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17], there is a fundamental issue of how to evaluate dynamical fluctuations. TURs usually involve fluctuations in fluxes or observables’ change [18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36], hence they require statistics more than single time points. However, in quantum mechanics, physical quantities at different times do not commute in the Heisenberg sense, which makes fluctuations elusive.
Conventionally, there are two approaches trying to characterize quantum fluctuations: one is based on the full counting statistics or continuous measurement [4, 5, 6, 7, 8, 9, 10, 11, 12, 13], which identifies jumps as exchanging charges with the environment or detections of a signal in an experiment [37]. However, it cannot explore fluctuations of more general observables that are not directly related to the jumps. The other approach utilizes the two-point measurement [14, 15, 16, 17], which, however, ends up with discarding coherence regarding the measured observable we are interested in, because it includes invasive measurement steps [38]. Hence, these methods can fail to capture all the quantum effects on thermodynamic trade-off relations.
@@
In this Letter, we derive a quantum TUR by quantifying the dynamical fluctuation of an observable with such a quasiprobability for the first time (see Fig. 1). We consider Markovian open quantum systems and show that the short-time variance of an observable’s change, as assessed by the Terletsky–Margenau–Hill quasiprobability [52, 53], provides a short-time TUR bound on the EPR for any state and observable. Moreover, on the basis of the non-classicality of quasiprobability, we elucidate criteria required for the anomalously large fluctuation absent in classical systems. Combined with our TUR, the criteria serve as a more fundamental, basis-independent condition for the dissipationless current [54, 55] than the abundance of the quantum coherence indicated in Ref. [54].
Setup.— We consider completely positive Markovian quantum dynamics, which are generally described by the quantum master equation [56, 57, 58]
@@
where $\rho$ is the density operator, $H$ the Hamiltonian, $L_{k}$ jump operators, $\cdot$$\cdot$] the commutator, and $\{\cdot,\cdot\}$ the anticommutator. We refer to $\mathcal{L}$ and $\mathcal{D}$ as the Liouvillian and the dissipator, respectively. Given the initial data $\rho_{0}$, Eq. (2) is formally solved by $\rho(t)=e^{\mathcal{L}t}\rho_{0}$. We indicate the adjoint of $\mathcal{L}$ and $\mathcal{D}$ by $\mathcal{L}^{\dagger}$ and $\mathcal{D}^{\dagger}$. For an observable (Hermitian operator) $X$, we describe the eigendecomposition as $X=\sum_{x}x\Pi_{x}$ with eigenvalues $x$ and projectors $\Pi_{x}$. Its expectation value at time $t$ is given by $\langle X\rangle_{t}=\mathop{\mathrm{tr}}\nolimits(X\rho(t))=\sum_{x}xp(x,t)$, where $p(x,t)=\mathop{\mathrm{tr}}\nolimits(\Pi_{x}\rho(t))$. We divide the time derivative into the Hamiltonian part $J_{X}^{\mathrm{H}}(t)=i\langle[H,X]\rangle_{t}$ and the dissipative part $J_{X}^{\mathrm{d}}(t)=\langle\mathcal{D}^{\dagger}(X)\rangle_{t}$ as $d_{t}\langle X\rangle_{t}=J_{X}^{\mathrm{H}}(t)+J_{X}^{\mathrm{d}}(t)$.
To discuss multi-time statistics of $X$ in the quantum dynamics beyond the single-time one $p(x,t)$, we introduce the quasiprobability [39, 40, 41]. In particular, we focus on the two-time statistics given by the Terletsky–Margenau–Hill (TMH) quasiprobability [52, 53], which is defined by
@@
for any state $\rho(t)$ and observable $X$. The numerator is the dissipative part of the time derivative of $\langle X\rangle_{t}$, so it quantifies the changing rate of $X$ due to the dissipative dynamics. On the other hand, the denominator, the short-time fluctuation, represents the dynamical fluctuation of $X$ as evaluated by the TMH quasiprobability. Thus, the inequality represents the universal trade-off between dissipation and fluctuations, i.e., the TUR [2, 3]. Specifically, it generalizes the so-called short-time TURs in classical systems [67, 68], which have been utilized in estimating the EPR in classical systems [69, 70, 68, 71, 72]. This connection will be clear if we assume that $\rho$ commutes with $X$; then, the numerator and denominator read $J_{X}^{\mathrm{d}}(t)=d_{t}\langle X\rangle_{t}$ and $m_{X}(t)=\sum_{x,y}(x-y)^{2}R_{yx}p(x,t)$, which appear in the short-time TUR presented in Ref. [68]. The TUR indicates that increasing the short-time fluctuation can lead to the reduction of the lower bound of dissipation. As will be discussed in detail later, the quasiprobabilistic perspective has the practical advantage of clarifying physical requirements for such an increase in $m_{X}$.
@@
Comparison with existing TURs.—We compare our TUR (10) with existing quantum TURs. First, we stress the difference from widely studied methods, the full counting statistics (FCS) approach [37] and the two-point measurement (TPM) approach [38]. In short, our result is complementary to those previous ones. While the TURs based on the FCS [6, 7, 8, 9, 10, 11, 12, 13] deal with current observables associated with the counting of jumps, our TUR is given by the fluctuations of intrinsic observables, which do not have to be connected to jumps. It is noteworthy that in classical Markov jump processes, an intrinsic observable’s change is entirely given by keeping track of jumps, because every physical quantity takes determined values in classical states [61]. However, quantum states linked by quantum jumps are not eigenstates of every intrinsic observable (they may not be even for the Hamiltonian; see Model A in [37]). In End Matter, we discuss a nontrivial intersection between the FCS and quasiprobabilities when $X,L_{k}]=w_{k}L_{k}$ holds for an observable $X$ with certain weights $w_{k}$.
The TPM method, adopted to quantify dynamical fluctuations in Refs. [14, 15, 16, 17], inevitably leads to decoherence in the initial density operator due to invasive initial measurement, despite having a clear experimental perspective. On the other hand, in our TUR, every initial coherence is incorporated owing to the employment of the quasiprobability, while maintaining experimental access by the interferometric method [40, 39, 62].
@@
Anomalous scaling via non-classicality of quasiprobability.— We next explore the ramifications of non-classicality on the thermodynamic constraint. As will be shown, the short-time fluctuation $m_{X}(t)$ can exhibit an anomalous scaling when $X$ is highly degenerate, which is forbidden in classical systems. This scaling makes the TUR bound small and allows for reducing entropy production beyond the classical limit. While the reduction has recently been attributed to quantum coherence in $\rho$ [54], we discuss that non-classical behaviors of the TMH quasiprobability are more fundamental.
First, we review the discussion in Ref. [54]. The authors derived an inequality similar to Eq. (10) for $X=H$, the Hamiltonian, where they used the coherence in $\rho$ regarding a specific eigenbasis of $H$ instead of the short-time fluctuation $m_{H}$. They showed that when $H$ is highly degenerate, dissipationless heat current can be realized, as $\dot{\Sigma}(\rho(t))=O(1)$ and $|d_{t}\langle H\rangle_{t}|=|J_{H}^{\mathrm{d}}(t)|=O(N)$, where $N$ is the number of degeneracy. In our terminology, this dissipationless current is enabled by $m_{H}(t)$ scaling at $O(N^{2})$, while they attributed it to the coherence’s $O(N)$ scaling.
To clarify what is crucial for the dissipationless current, we formulate the problem in a general way. We assume that the eigenvalues of $X$ are $N$-fold degenerate and examine conditions on $\rho(t)$ for $m_{X}(t)$ to scale more rapidly than the classical limitation $m_{X}(t)=O(N)$ (detailed below). We specifically refer to the growth of the short-time fluctuation faster than $O(N)$ as anomalous scaling. When an eigenbasis for eigenvalue $x_{s}$ is given as $\{|s,j\rangle\}_{j=1}^{N}$, $X$ is expanded as $X=\sum_{s,j}x_{s}\Pi_{s,j}$ with $\Pi_{s,j}=|s,j\rangle\langle s,j|$. We can define the flux from $(s,j)$ to $(s^{\prime},j^{\prime})$ as $T_{s^{\prime}j^{\prime}sj}(\rho)=\frac{1}{2}\mathop{\mathrm{tr}}\nolimits(\{\mathcal{L}^{\dagger}\Pi_{s^{\prime},j^{\prime}},\Pi_{s,j}\}\rho)$ and the integrated fluxes as
$\mathcal{T}_{s^{\prime}s}(\rho)$ $\coloneqq\begin{cases}\sum_{j,j^{\prime}}T_{s^{\prime}j^{\prime}sj}(\rho)&(s\neq s^{\prime}),\\ \sum_{j,j^{\prime}(\neq j)}T_{sj^{\prime}sj}(\rho)&(s=s^{\prime}).\end{cases}$ (15)
While $T_{s^{\prime}j^{\prime}sj}$ and $\mathcal{T}_{ss}$ are basis-dependent, $\mathcal{T}_{s^{\prime}s}$ are not if $s\neq s^{\prime}$. We further define the average escape rate $\bar{\lambda}(\rho)=-\sum_{s,j}T_{sjsj}(\rho)$. They satisfy the equality $\sum_{s,s^{\prime}}\mathcal{T}_{s^{\prime}s}(\rho)-\bar{\lambda}(\rho)=0$. Because the time dependence is not relevant hereafter, we write $m_{X}(\rho)$ to indicate $m_{X}(t)$ when $\rho(t)=\rho$. The short-time fluctuation is given solely by the integrated fluxes as $m_{X}(\rho)=\sum_{s,s^{\prime}}(x_{s}-x_{s^{\prime}})^{2}\mathcal{T}_{s^{\prime}s}(\rho)$.
We first demonstrate $m_{X}(\rho)=O(N)$ in the classical cases. To this end, we define an eigenbasis to be classical if it satisfies that for each $(s,j,s^{\prime},j^{\prime})$, the number of jumps such that $|\langle s^{\prime},j^{\prime}|L_{k}|s,j\rangle|\neq 0$ is $O(1)$, and that $|\langle s^{\prime},j^{\prime}|L_{k}|s,j\rangle|=O(1)$. For a classical eigenbasis, we find $R_{s^{\prime}j^{\prime}sj}=\sum_{k}|\langle s^{\prime},j^{\prime}|L_{k}|s,j\rangle|^{2}=O(1)$ and $R_{sjsj}=\sum_{s^{\prime},j^{\prime}(\neq s,j)}\sum_{k}|\langle s^{\prime},j^{\prime}|L_{k}|s,j\rangle|^{2}=O(N)$. Therefore, if $\rho$ has no coherence with respect to the classical eigenbasis (i.e., $\rho=\sum_{s,j}p_{sj}\Pi_{s,j}$ holds), we obtain the scaling $\bar{\lambda}(\rho)=-\sum_{s,j}R_{sjsj}p_{sj}=O(N)$ because $p_{sj}\geq 0$ and $\sum_{s,j}p_{sj}=1$. This scaling is intuitive as the target of escape grows by at most $O(N)$. Because the absence of coherence in $\rho$ also leads to the positivity of $T_{s^{\prime}j^{\prime}sj}(\rho)$, we can conclude $m_{X}(\rho)=O(N)$ from
$m_{X}(\rho)=\sum_{s,s^{\prime}}(x_{s}-x_{s}^{\prime})^{2}\mathcal{T}_{s^{\prime}s}(\rho)$
$\leq\max_{s,s^{\prime}|\mathcal{T}_{s^{\prime}s}(\rho)\neq 0}(x_{s}-x_{s^{\prime}})^{2}\sum_{s,s^{\prime}}\mathcal{T}_{s^{\prime}s}(\rho)$
$=\max_{s,s^{\prime}|\mathcal{T}_{s^{\prime}s}(\rho)\neq 0}(x_{s}-x_{s^{\prime}})^{2}\bar{\lambda}(\rho)=O(N).$
Here, we have assumed that $\mathcal{T}_{s^{\prime}s}(\rho)\neq 0$ implies $|x_{s}-x_{s^{\prime}}|=O(1)$, which is true if, for example, there is a degeneracy-independent threshold $\Lambda$ such that $|x_{s}-x_{s^{\prime}}|>\Lambda$ implies $\langle s^{\prime},j^{\prime}|H|s,j\rangle=0$ and $\langle s^{\prime},j^{\prime}|L_{k}|s,j\rangle=0$ (namely, transitions leading to a drastic change in $X$ are forbidden).
The above discussion shows that if $\rho$ has no coherence regarding the classical eigenbases, $m_{X}(\rho)$ grows by at most $O(N)$. On the other hand, if we adopt a non-classical eigenbasis, the coherence tells us nothing. For example, consider $\rho$ expanded as $\rho=\sum_{s}p_{s}|s,*\rangle\langle s,*|$ with $|s,*\rangle$ such that $X|s,*\rangle=x_{s}|s,*\rangle$ and $|\langle s^{\prime},*^{\prime}|L_{k}|s,*\rangle|=O(N^{\alpha})$ with $\alpha>1/2$. Then, $\bar{\lambda}$ grows by $O(N^{2\alpha})>O(N)$ and an anomalous scaling can occur even though $\rho$ has no coherence. We will exemplify this drawback of coherence-based consideration through an example later.
@@
The next statement is our second main result, which provides another viewpoint to examine the anomalous scaling: if neither of the following conditions is satisfied for some eigenbasis, $m_{X}(\rho)$ does not exhibit anomalous scaling:
@@
$q((s^{\prime},j^{\prime}),t+\Delta t;(s,j),t)$
@@
Note that the negativity of the quasiprobability also indicates the genuine quantumness called contextuality [43, 44]. The other one (Q2) is also non-classical because it means the chance of escape grows more rapidly than the number of evacuation targets increases. The statement shows that either of these quantumness conditions has to hold in any basis for the anomalous scaling to occur. Such a property stands in contrast to the existence of coherence, which does not need to hold in every basis. We depict their relationship with anomalous scaling in Fig. 2.
Figure 2: Hierarchy of conditions. Condition (Q1) or (Q2) is always necessary for $m_{X}(\rho)$ to scale anomalously, as is prohibited in classical systems. (a) For classical eigenbases, these conditions are sufficient for the state to have coherence. Thus, this time, coherence is necessary for the anomalous scaling. Density matrix $\rho^{+}$, discussed in Example, has a large coherence regarding a classical basis to exhibit anomalous scaling. (b) If we take a non-classical eigenbasis, coherence regarding that basis loses its connection to the scaling ($\rho^{+}$ may become “incoherent”). On the other hand, our two conditions remain relevant. In both cases, there can be a state $\rho^{-}$ that has as much coherence as $\rho^{+}$ but does not satisfy (Q1) or (Q2) and thus does not show anomalous scaling.
@@
Let us prove the claim. First, because $\sum_{s,s^{\prime}}\mathcal{T}_{s^{\prime}s}(\rho)=\bar{\lambda}(\rho)$, the negation of (Q2) implies that $\sum_{s,s^{\prime}}\mathcal{T}_{s^{\prime}s}(\rho)$ is $O(N)$. Then, split the summation as $\sum_{s,s^{\prime}}\mathcal{T}_{s^{\prime}s}(\rho)=\mathcal{T}^{+}-\mathcal{T}^{-}$ with $\mathcal{T}^{+}=\sum_{s,s^{\prime}|\mathcal{T}_{s^{\prime}s}(\rho)\geq 0}\mathcal{T}_{s^{\prime}s}(\rho)$ and $\mathcal{T}^{-}=\sum_{s,s^{\prime}|\mathcal{T}_{s^{\prime}s}(\rho)<0}|\mathcal{T}_{s^{\prime}s}(\rho)|$. Due to the negation of (Q1), the negative part $\mathcal{T}^{-}$ is at most $O(N)$. Therefore, the positive part $\mathcal{T}^{+}$ can also be at most $O(N)$. Finally, $m_{X}(\rho)$ is bounded as

### v2

@@ Abstract
We derive a quantum extension of the thermodynamic uncertainty relation where dynamical fluctuations are quantified by the Terletsky–Margenau–Hill quasiprobability, a quantum generalization of the classical joint probability. The obtained inequality plays a complementary role to existing quantum thermodynamic uncertainty relations, focusing on observables’ change rather than exchange of charges through jumps and respecting initial coherence. Quasiprobabilities show anomalous behaviors that are forbidden in classical systems, such as negativity; we reveal that negativity or a non-classically enhanced escape rate is necessary to increase an output-to-dissipation ratio beyond classical limitations and show that the requirements are basis-independent and stronger than quantum coherence. To illustrate these statements, we employ a model that can exhibit a dissipationless heat current, which would be prohibited in classical systems; we construct a state that has much coherence but does not lead to a dissipationless current due to the absence of anomalous behaviors in quasiprobabilities.
@@
While extending the TUR to the quantum regime is a crucial problem gathering much attention recently [4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20], there is a fundamental issue of how to evaluate dynamical fluctuations. TURs usually involve fluctuations in fluxes or observables’ change [21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39], hence they require statistics more than single time points. However, in quantum mechanics, physical quantities at different times do not commute in the Heisenberg sense, which makes fluctuations elusive.
Conventionally, there are two approaches trying to characterize quantum fluctuations: one is based on the full counting statistics or continuous measurement [4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14], which identifies jumps as exchanging charges with the environment or detections of a signal in an experiment [40]. However, it cannot explore fluctuations of more general observables that are not directly related to the jumps. The other approach utilizes the two-point measurement [16, 17, 18, 19], which, however, ends up with discarding coherence regarding the measured observable we are interested in, because it includes invasive measurement steps [41]. Hence, these methods can fail to capture all the quantum effects on thermodynamic trade-off relations.
@@
In this Letter, we derive a quantum TUR by quantifying the dynamical fluctuation of an observable with such a quasiprobability for the first time (see Fig. 1). We consider Markovian open quantum systems and show that the short-time variance of an observable’s change, as assessed by the Terletsky–Margenau–Hill quasiprobability [55, 56], provides a short-time TUR bound on the EPR for any state and observable. Moreover, on the basis of the non-classicality of quasiprobability, we elucidate criteria required for the anomalously large fluctuation absent in classical systems. Combined with our TUR, the criteria show that a large amount of heat flux with finite dissipation as reported in Refs. [57, 58] can happen only when the quasiprobability exhibits negative values or a non-classically enhanced escape rate. We further compare our quasiprobability-based conditions with the coherence-based ones discussed in Ref. [57].
Setup.— We consider completely positive Markovian quantum dynamics, which are generally described by the Gorini–Kossakowski–Sudarshan–Lindblad (GKSL) equation [59, 60, 61]
@@
where $\rho$ is the density operator, $H$ the Hamiltonian, $L_{k}$ jump operators, $\cdot$$\cdot$] the commutator, and $\{\cdot,\cdot\}$ the anticommutator. We refer to $\mathcal{L}$ and $\mathcal{D}$ as the Liouvillian and the dissipator, respectively. While we assume the Hamiltonian and jump operators are time-independent, our results are easily extended to time-dependent cases. Given the initial data $\rho_{0}$, Eq. (2) is formally solved by $\rho(t)=e^{\mathcal{L}t}\rho_{0}$. We indicate the adjoint of $\mathcal{L}$ and $\mathcal{D}$ by $\mathcal{L}^{\dagger}$ and $\mathcal{D}^{\dagger}$. For an observable (Hermitian operator) $X$, we describe the eigendecomposition as $X=\sum_{x}x\Pi_{x}$ with eigenvalues $x$ and projectors $\Pi_{x}$. Its expectation value at time $t$ is given by $\langle X\rangle_{t}=\mathop{\mathrm{tr}}\nolimits(X\rho(t))=\sum_{x}xp(x,t)$, where $p(x,t)=\mathop{\mathrm{tr}}\nolimits(\Pi_{x}\rho(t))$. We divide the time derivative into the Hamiltonian part $J_{X}^{\mathrm{H}}(t)=i\langle[H,X]\rangle_{t}$ and the dissipative part $J_{X}^{\mathrm{d}}(t)=\langle\mathcal{D}^{\dagger}(X)\rangle_{t}$ as $d_{t}\langle X\rangle_{t}=J_{X}^{\mathrm{H}}(t)+J_{X}^{\mathrm{d}}(t)$.
To discuss multi-time statistics of $X$ in the quantum dynamics beyond the single-time one $p(x,t)$, we introduce the quasiprobability [42, 43, 44]. In particular, we focus on the two-time statistics given by the Terletsky–Margenau–Hill (TMH) quasiprobability [55, 56, 62], which is defined by
@@
for any state $\rho(t)$ and observable $X$. The numerator is the dissipative part of the time derivative of $\langle X\rangle_{t}$, so it quantifies the changing rate of $X$ due to the dissipative dynamics. On the other hand, the denominator, the short-time fluctuation, represents the dynamical fluctuation of $X$ as evaluated by the TMH quasiprobability. Thus, the inequality represents the universal trade-off between dissipation and fluctuations, i.e., the TUR [2, 3]. Specifically, it generalizes the so-called short-time TUR in classical systems [71] (also known as the entropic bound [72, 36]), which have been utilized in estimating the EPR [73, 74, 71, 75, 76] and deriving the power-efficiency trade-off [77, 36]. This connection will become clear if we assume that $\rho$ commutes with $X$; then, the numerator and denominator read $J_{X}^{\mathrm{d}}(t)=d_{t}\langle X\rangle_{t}$ and $m_{X}(t)=\sum_{x,y}(x-y)^{2}R_{yx}p(x,t)$, which appear in the short-time TUR presented in Ref. [71]. As discussed in Ref. [36], the classical short-time TUR (entropic bound) holds without modification when the system is under a time-dependent protocol or has odd variables  [^1] Note that Ref. [36] presents the bound in the time-integrated form, but it is equivalent to the instantaneous form., unlike the finite-time versions [29, 33]; our TUR also applies to time-dependent systems, and we expect that it holds even if the system’s dynamics is not time-reversal symmetric.
@@
Comparison with existing TURs.—We compare our TUR (10) with existing quantum TURs. First, we stress the difference from widely studied methods, the full counting statistics (FCS) approach [40] and the two-point measurement (TPM) approach [41]. In short, our result is complementary to those previous ones. While the TURs based on the FCS [6, 7, 8, 9, 10, 11, 13, 14, 15] deal with current observables associated with the counting of jumps, our TUR is given by the fluctuations of intrinsic observables (i.e., Hermitian operators on the system’s Hilbert space), which do not have to be connected to jumps. It is noteworthy that in classical Markov jump processes, an intrinsic observable’s change is entirely given by keeping track of jumps, because every physical quantity takes determined values in classical states [65]. However, quantum states linked by quantum jumps are not eigenstates of every intrinsic observable (they may not be even for the Hamiltonian; see Model A in [40]). In End Matter, we discuss a nontrivial intersection between the FCS and quasiprobabilities when $X,L_{k}]=w_{k}L_{k}$ holds for an observable $X$ with certain weights $w_{k}$.
The TPM method, adopted to quantify dynamical fluctuations in Refs. [16, 17, 18, 19], inevitably leads to decoherence in the initial density operator due to invasive initial measurement, despite having a clear experimental perspective. Thus, derived TURs fail to capture initial coherence’s effects. On the other hand, in our TUR, every initial coherence is incorporated owing to the employment of the quasiprobability, while maintaining experimental access by the interferometric method [43, 42, 66].
@@
Anomalous scaling via non-classicality of quasiprobability.— When rewritten as
$Q_{X}(\rho)\coloneqq\frac{|J_{X}^{\mathrm{d}}(\rho)|^{2}}{\dot{\Sigma}(\rho)}\leq\frac{1}{2}m_{X}(\rho),$ (15)
the TUR shows that the output-to-dissipation ratio $Q_{X}(\rho)$ is upper bounded by the short-time fluctuation (hereafter, we write $J_{X}^{\mathrm{d}}(\rho)$ and $m_{X}(\rho)$ to indicate $J_{X}^{\mathrm{d}}(t)$ and $m_{X}(t)$ when $\rho(t)=\rho$). When $X$ is extensive, it is expected that $Q_{X}(\rho)$ can grow at most at the order of system size. Intriguingly, in Ref. [57], the authors show that in a system with an $N$-fold degenerate Hamiltonian $H$, we can let $J_{H}(\rho)$ grow at $O(N)$, keeping $\dot{\Sigma}(\rho)=O(1)$, which means $Q_{H}(\rho)=O(N^{2})$, violating the classical limit of $O(N)$. They call such a phenomenon dissipationless current. They also discuss that for a certain class of degenerate eigenbases, the dissipationless current can be attributed to a large amount of quantum coherence in $\rho$.
Owing to the quasiprobability TUR (10), which connects dissipation and quasiprobabilities, we can characterize this non-classical phenomenon in an eigenbasis-free way. We investigate the scaling of $m_{X}(\rho)$ for an $N$-fold degenerate observable $X$ instead of $Q_{X}(\rho)$ because its fast scaling is a necessary condition from Eq. (15). To understand the effect of degeneracy on the scaling, we assume that the variety of $X$’s eigenvalue and its operator norm are $O(1)$ and that the degree of degeneracy is independent of the eigenvalue; thus, we assume the form $X=\sum_{s}\sum_{j=1}^{N}x_{s}\Pi_{s,j}$ with $\Pi_{s,j}=|s,j\rangle\langle s,j|$ and $X|s,j\rangle=x_{s}|s,j\rangle$. This includes the Hamiltonian considered in Ref. [57]. We call the $O(N^{\alpha})$ scaling of $m_{X}(\rho)$ with $\alpha>1$ anomalous scaling.
Before stating the result, we introduce a couple of quantities. For an eigenbasis, the flux from $(s,j)$ to $(s^{\prime},j^{\prime})$ and the average escape rate are respectively defined as $T_{s^{\prime}j^{\prime}sj}(\rho)=\frac{1}{2}\mathop{\mathrm{tr}}\nolimits(\{\mathcal{L}^{\dagger}\Pi_{s^{\prime},j^{\prime}},\Pi_{s,j}\}\rho)$ and $\bar{\lambda}(\rho)=-\sum_{s,j}T_{sjsj}(\rho)$. We further define the integrated fluxes $\mathcal{T}_{s^{\prime}s}(\rho)$ as
$\mathcal{T}_{s^{\prime}s}(\rho)$ $\coloneqq\begin{cases}\sum_{j,j^{\prime}}T_{s^{\prime}j^{\prime}sj}(\rho)&(s\neq s^{\prime})\\ \sum_{j,j^{\prime}(\neq j)}T_{sj^{\prime}sj}(\rho)&(s=s^{\prime})\end{cases}.$ (16)
The equality $\sum_{s,s^{\prime}}\mathcal{T}_{s^{\prime}s}(\rho)-\bar{\lambda}(\rho)=0$ follows from the property $\sum_{s^{\prime},j^{\prime}}T_{s^{\prime}j^{\prime}sj}(\rho)=0$. The short-time fluctuation is given solely by the integrated fluxes as $m_{X}(\rho)=\sum_{s,s^{\prime}}(x_{s}-x_{s^{\prime}})^{2}\mathcal{T}_{s^{\prime}s}(\rho)$.
@@
The next statement is our second main result: if neither of the following conditions is satisfied for some eigenbasis, $m_{X}(\rho)$ does not exhibit anomalous scaling:
@@
$q\big((s^{\prime},j^{\prime}),t+\Delta t;(s,j),t\big)$
@@
Note that the negativity of the quasiprobability also indicates the genuine quantumness called contextuality [46, 47]. The other one (Q2) is also non-classical because it means the chance of escape grows more rapidly than the number of evacuation targets increases. The statement shows that either of these quantumness conditions has to hold in any basis for the anomalous scaling to occur.
When $\rho$ commutes with $X$, the quasiprobability conditions necessitate non-classicality at the level of basis components for every simultaneous eigenbasis. Now, we define an eigenbasis $\{|s,j\rangle\}$ to be classical if the following conditions are satisfied for all $(s,j)$ and $(s^{\prime},j^{\prime})(\neq(s,j))$: as $N$ increases, (1) the jump rate does not grow ($|\langle s^{\prime},j^{\prime}|L_{k}|s,j\rangle|=O(1)$), and (2) the number of jumps that mediate the states does not grow ($\{k\colon|\langle s^{\prime},j^{\prime}|L_{k}|s,j\rangle|>0\}=O(1)$). Then, if a simultaneous eigenbasis $\{|s,j\rangle\}$ of $\rho$ and $X$ is classical, we can prove that the anomalous scaling does not occur. First observe that $R_{sjsj}=\sum_{s^{\prime},j^{\prime}(\neq s,j)}\sum_{k}|\langle s^{\prime},j^{\prime}|L_{k}|s,j\rangle|^{2}=O(N)$, and thus $\bar{\lambda}(\rho)=-\sum_{s,j}R_{sjsj}p_{sj}=O(N)$, where $p_{s,j}=\mathop{\mathrm{tr}}\nolimits(\rho\Pi_{s,j})$. Besides, we have $T_{s^{\prime}j^{\prime}sj}(\rho)=\sum_{k}|\langle s^{\prime},j^{\prime}|L_{k}|s,j\rangle|^{2}p_{s,j}\geq 0$ since $\rho$ is diagonal with respect to $\{|s,j\rangle\}$. Thus, $m_{X}(\rho)$ can grow at most at $O(N)$ because both (Q1) and (Q2) are violated by the eigenbasis. This is in accordance with the result of Ref. [57], where dissipationless current is proved not to occur when $\rho$ has no coherence with respect to this type of eigenbasis.
In addition, we can show that neither (Q1) nor (Q2) is realized unless $\rho$ has significant (more than $O(1)$) coherence [80]; hence, our quasiprobability-based conditions are tighter than having divergent coherence, as depicted in Fig. 2(a). Moreover, given at the quasiprobability level, the criteria apply to any eigenbasis (see Fig. 2(b))  [^2] When a non-classical eigenbasis is used, the corresponding classical Markov jump process has transition rates that scale as $O(N^{\alpha})$ with some positive exponent $\alpha$. However, this scaling is typically not possible in realistic classical systems [77], and therefore the classicality of the eigenbasis is implicitly assumed in Ref. [57].. The relationship is illustrated later through an example.
Figure 2: Hierarchy of conditions. Condition (Q1) or (Q2) is always required for $m_{X}(\rho)$ to scale anomalously, which is necessary for the output-to-dissipation ratio $Q_{X}(\rho)$ to grow faster than the classical limit. (a) For classical eigenbases, our necessary conditions are tighter than the condition that the state has more than $O(1)$ coherence. (b) If we take a non-classical eigenbasis, coherence regarding that basis loses its connection to the scaling. In contrast, our two conditions remain relevant. In the example, we examine two characteristic cases, $\rho^{+}$ and $\rho^{-}$.
@@
Let us prove the main claim that either (Q1) or (Q2) is necessary for anomalous scaling. First, because $\sum_{s,s^{\prime}}\mathcal{T}_{s^{\prime}s}(\rho)=\bar{\lambda}(\rho)$, the negation of (Q2) implies that $\sum_{s,s^{\prime}}\mathcal{T}_{s^{\prime}s}(\rho)$ is $O(N)$. Then, split the summation as $\sum_{s,s^{\prime}}\mathcal{T}_{s^{\prime}s}(\rho)=\mathcal{T}^{+}-\mathcal{T}^{-}$ with $\mathcal{T}^{+}=\sum_{s,s^{\prime}|\mathcal{T}_{s^{\prime}s}(\rho)\geq 0}\mathcal{T}_{s^{\prime}s}(\rho)$ and $\mathcal{T}^{-}=\sum_{s,s^{\prime}|\mathcal{T}_{s^{\prime}s}(\rho)<0}|\mathcal{T}_{s^{\prime}s}(\rho)|$. Due to the negation of (Q1), the negative part $\mathcal{T}^{-}$ is at most $O(N)$. Therefore, the positive part $\mathcal{T}^{+}$ can also be at most $O(N)$. Finally, $m_{X}(\rho)$ is bounded as